Three girls were stringing beads. There were a total of 195 beads in all and they were shared in a ratio of 1/2:1/3:1/4. How many beads did each of the girls receive?

Answer by KMST(5328)

ONE WAY TO LOOK AT IT:
Multiplying the proportions 1/2:1/3:1/4 times 12 we get
12/2:12/3:12/4 or 6:4:3 .
That means that for every 6 + 4 + 3 = 13 beads, they got 6, 4, and 3 beads respectively.
195/13 = 15 , so we can distribute 13 beads 15 times
with one girl receiving 15*6 = 90 beads,
the other one getting 15*4 = 60 beads,
and the third one getting 15*3 = 45 beads.

ANOTHER WAY:
They got (1/2)x, (1/3)x, and (1/4)x respectively for some number x, and
(1/2)x + (1/3)x + (1/4)x = 195
(1/2+1/3+1/4)x = 195
(6/12+4/12+3/12)x = 195
(13/12)x = 195
x = 195(12/13)
x = 195*12/13
x = 180
So they got
(1/2)180 = 90,
(1/3)180 = 60, and
(1/4)180 = 45 respectively.